OpenStudy (anonymous):

5/3x+1=3/3x-1 what does x equal

OpenStudy (anonymous):

3x = 3x-2?

OpenStudy (anonymous):

\[\frac{5}{3x}+1=\frac{3}{3x-1}\] ?

OpenStudy (anonymous):

its 5 over 3x+1

OpenStudy (anonymous):

\[\frac{5}{3x+1}=\frac{3}{3x-1}\] Cross Multiply 5(3x-1)=3(3x+1) 15x-5=9x+3

OpenStudy (anonymous):

what do i do after that?

OpenStudy (anonymous):

6x=8